Order: 450mL of 30% Osmolite solution q6h through an NG tube. How much Osmolite and water should be mixed to equal 450mL?
Answer #1

Ans) Medication Order: 450 mL of a 30% OsCal solution through a NG tube.

Next, Calculate volume of water.

When ordered at less than full strength, enteral feedings are ordered in percent values. In this case, the nurse must calculate the amount of feeding solution and the amount of water needed. This can be done using the formula method, the ratio and proportion method, a fractional equation, or the dimensional analysis method.

Desired percentage (D%) = 30%

Percent on hand (H%) = 100%

Desired total Volume (V) = 450ml

Aao, 30/100 × 450 = 135ml

Calculating the volume of Medication in ml:

13500/100 = 135ml of OsCal.

Subtract the known medication amount from the total volume to get the volume of water.
Total volume - known volume of drug = Volume of water
450 - 135 = 315ml